# Evaluating Human Alignment and Model Faithfulness of LLM Rationale

### v1

@@ Appendix A Appendix
A.1 Top-k Faithfulness
As previously noted, the number of masked words significantly impacts the flip rate. To explore this further, we conducted a Top-$k$ experiment, masking $k=1,2,3,4,5,10$ words and calculating the flip rate (see Table 8). The results consistent with Table 5 indicate that on average, attribution-based methods surpass prompting in terms of faithfulness.
Figure 4: Token Importance before and after fine-tuning Llama-2-7B based on the InputXGradient explainability method. The predicted word by the model is shown in green and its attributions to previous words are shown in red. The attributions before fine-tuning are more skewed (Fisher-Pearson coefficient of skewness over all the dataset: 3.37±0.31), and become less skewed after fine-tuning (1.42±0.36).
Figure 5: Token Importance before and after fine-tuning Llama-2-7B based on the Input×Gradient explainability method. The predicted/true label is shown in green, with its attributions to input words in red. The human rationale for the examples are ["dog", "cat], ["twins"], and ["grazes", "touching"]. The fine-tuned model identified these words solely through training on classification data, without any rationale data.
@@ A.1 Top-k Faithfulness
Table 6: The details of the models we used in this work.
@@
Learning Rate E-SNLI 1e-05
Learning Rate MedicalBios 5e-06
Num Epochs 5
@@
Table 7: The hyperparameters used for fine-tuning the models using LoRA.
Model Mistral-7B FT Instruct-v0.2 Llama-2-7b FT chat Llama-3-8B FT Instruct
Top k 1 2 3 4 5 10 Avg 1 2 3 4 5 10 Avg 1 2 3 4 5 10 Avg
Dataset Method
E-SNLI Attention 41.0 36.7 48.0 47.7 54.0 62.3 48.3 40.0 50.0 63.7 66.3 69.3 71.0 60.1 31.0 35.3 41.7 49.7 56.0 69.7 47.2
Saliency 40.0 46.0 49.3 52.3 54.7 63.0 50.9 36.7 45.0 54.3 58.0 59.7 68.3 53.7 30.3 37.7 50.0 56.3 63.7 77.0 52.5
Input×Gradient 34.0 39.0 44.0 46.7 50.0 57.7 45.2 37.0 42.3 50.7 60.3 60.7 70.0 53.5 33.0 38.0 54.3 59.0 61.0 77.3 53.8
Prompting 31.3 30.0 41.3 40.7 47.0 55.3 40.9 35.3 44.3 54.7 60.7 64.3 64.0 53.9 29.3 28.7 47.0 48.7 57.7 76.7 48.0
MedicalBios Attention 7.1 10.6 15.0 15.0 19.5 31.0 16.4 5.3 8.0 7.1 8.8 9.7 13.3 8.7 0.9 3.6 3.6 5.4 6.2 15.2 5.8
Saliency 5.3 12.4 14.2 17.7 20.4 31.0 16.8 14.2 18.6 18.6 15.9 16.8 18.6 17.1 7.1 11.6 13.4 17.0 18.8 18.8 14.4
Input×Gradient 8.8 16.8 15.0 14.2 17.7 25.7 16.4 15.9 17.7 19.5 15.0 18.6 19.5 17.7 8.0 12.5 14.3 16.1 17.0 22.3 15.0
Prompting 7.1 10.6 16.8 16.8 17.7 28.3 16.2 0.9 9.7 14.2 19.5 17.7 19.5 13.6 6.2 7.1 12.5 19.6 16.1 27.7 14.9
[…]
Model Mistral-7B FT Instruct-v0.2 Llama-2-7b FT chat Llama-3-8B FT Instruct
Top k 1 2 3 4 5 10 1 2 3 4 5 10 1 2 3 4 5 10
Dataset Method
E-SNLI Attention 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0
Saliency 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0
Input×Gradient 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0
Prompting 1.0 1.8 2.9 3.8 4.7 8.2 0.9 1.9 2.9 3.8 4.8 7.4 0.9 1.8 2.8 3.8 4.8 9.1
MedicalBios Attention 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0
Saliency 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0
Input×Gradient 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0 1.0 2.0 3.0 4.0 5.0 10.0
Prompting 1.0 1.8 2.9 3.9 4.9 9.9 0.2 1.6 2.7 3.7 4.7 8.5 0.8 1.7 2.7 3.7 4.5 7.7
Table 9: The average number of masked words in each experiment of Table 8. Prompting can have fewer masked words as a result of generating words out of the input sentence or not following the instruction on how many words it should generate (We limit the masks if the model generates more than $k$ words).
Model Mistral-7B Instruct-v0.2 Llama-2-7b chat Llama-3-8B Instruct GPT-3.5 Turbo 1106
Dataset E-SNLI MedicalBios E-SNLI MedicalBios E-SNLI MedicalBios E-SNLI MedicalBios
Masking Part Method Selection
Attribution-based
Input Saliency Top-Ratio 3.00 50.00 2.36 31.82 29.33 36.11 - -
Input×Gradient Top-Ratio 2.33 44.32 2.02 29.09 26.67 40.74 - -
Input&Instr. Saliency Top-Ratio 97.00 71.59 100.00 90.91 86.00 72.22 - -
Input×Gradient Top-Ratio 97.33 63.64 100.00 93.64 86.33 74.07 - -
Prompting-based
Input Normal Prompt Unbound 3.33 59.09 1.01 30.91 38.67 41.67 77.00 41.82
Short Prompt Unbound 3.67 59.09 1.68 35.45 44.00 47.22 73.67 37.27
Normal Prompt Top-Ratio 1.67 28.41 1.68 33.64 24.00 45.37 41.33 24.55
Short Prompt Top-Ratio 2.33 30.68 1.01 26.36 22.67 46.30 44.33 20.91
Input&Instr. Extended Prompt Unbound 68.67 61.36 8.75 28.18 58.33 45.37 99.67 50.00
Baselines
Input Human Human 4.00 60.23 2.02 38.18 31.67 43.12 55.67 36.36
Everything Everything 2.33 97.73 0.34 60.00 69.67 70.64 89.00 74.55
[…]
Model Mistral-7B FT Instruct-v0.2 Llama-2-7b FT chat Llama-3-8B FT Instruct
Dataset E-SNLI MedicalBios E-SNLI MedicalBios E-SNLI MedicalBios
Selection Method
Unbound Normal Prompt 64.00 34.51 67.00 17.70 64.33 18.75
Short Prompt 53.00 37.17 67.33 21.24 72.00 25.00
Top-Ratio Attention 50.67 28.32 70.33 15.04 53.67 9.82
Saliency 52.33 30.09 62.67 24.78 66.00 23.21
Input×Gradient 47.67 25.66 62.33 26.55 62.67 25.89
Normal Prompt 54.67 32.74 63.00 23.01 61.67 24.11
Short Prompt 52.67 35.40 64.00 25.66 59.00 25.00
Top-Var Attention 47.67 23.01 67.33 15.93 53.67 6.25
Saliency 50.00 19.47 58.67 26.55 60.00 17.86
Input×Gradient 44.33 20.35 58.00 24.78 61.67 19.64
Normal Prompt 51.33 30.97 64.67 19.47 61.33 21.43
Short Prompt 50.67 30.09 60.67 20.35 59.00 20.54
Baselines
TOP-VAR Human 49.67 21.24 60.00 24.78 62.00 24.11
Everything Everything 66.33 84.07 63.00 75.22 77.33 74.11
[…]
Figure 6: The human alignment (F1) and model faithfulness flip rate, evaluated at the intersection of correctly predicted examples (important for alignment) and the samples that were predicted within the output space of classes prior to fine-tuning (important for faithfulness). Additionally, the number of masks is constrained to the Top-Var selection.
@@
NORMAL PROMPT UNBOUND You will be provided a short premise and a hypothesis, each delimited with triple backticks.
Your task is to identify the most important key words in the premise and hypothesis that are most relevant to understanding the relationship between the premise and the hypothesis given a label (entailment/neutral/contradiction).
• These key words should be selected to capture the main idea or reasoning behind why the given label indicates the relationship between the premise and the hypothesis.
• The key words should be copied verbatim from the text.
• Write the key words as a pipe-separated (|) list of words with spaces.
• Example output: <word1> | <word2> | <word3>
• Make sure to only output keyword tokens and do not explain the answer.
Premise: “‘ {premise} “‘
@@
SHORT PROMPT UNBOUND Identify the most important key words from the text that are most relevant to understanding the relationship between the premise and the hypothesis.
Write the words as a pipe-separated (|) list of words with spaces and do not explain the answer.
Premise: “‘ {premise} “‘
@@
EXTENDED PROMPT UNBOUND Consider the following prompt: " Determine the inference relation between two (short, ordered) texts as entailment, if the premise entails the hypothesis, contradiction, if the hypothesis contradicts the premise, or neutral, if neither entailment nor contradiction hold.
Give your answer as "entailment", "neutral" or "contradiction" and nothing else.
@@
Label: "
Identify the most important single key words from the whole prompt (Including instruction and sentence) that are most relevant to understanding the relationship between the premise and the hypothesis as {label}.
Just write the words as a pipe-separated (|) list of words with spaces and do not explain the answer.
NORMAL PROMPT TOP-VAR / TOP-RATIO You will be provided a short premise and a hypothesis, each delimited with triple backticks.
Your task is to identify the top {k} most important key words in the premise and hypothesis that are most relevant to understanding the relationship between the premise and the hypothesis given a label (entailment/neutral/contradiction).
• These key words should be selected to capture the main idea or reasoning behind why the given label indicates the relationship between the premise and the hypothesis.
• The key words should be copied verbatim from the text.
• Write the key words as a pipe-separated (|) list of words with spaces.
• Example output: <word1> | <word2> | <word3>
• Make sure to only output top {k} keyword tokens and do not explain the answer.
Premise: “‘ {premise} “‘
Hypothesis: “‘ {hypothesis} “‘
@@
SHORT PROMPT TOP-VAR / TOP-RATIO Identify the top {k} most important key words from the text that are most relevant to understanding the relationship between the premise and the hypothesis.
Write the top {k} words as a pipe-separated (|) list of words with spaces and do not explain the answer.
Premise: “‘ {premise} “‘
Hypothesis: “‘ {hypothesis} “‘
@@
ATTRIBUTION-BASED TOP-VAR / TOP-RATIO
@@
Determine the inference relation between two (short, ordered) texts as entailment, if the premise entails the hypothesis, contradiction, if the hypothesis contradicts the premise, or neutral, if neither entailment nor contradiction hold.
Give your answer as "entailment", "neutral" or "contradiction" and nothing else.
Premise and Hypothesis: “‘ {premise} {hypothesis} “‘
Label: {predicted_class}
CLASSIFICATION Determine the inference relation between two (short, ordered) texts as entailment, if the premise entails the hypothesis, contradiction, if the hypothesis contradicts the premise, or neutral, if neither entailment nor contradiction hold.
Give your answer as "entailment", "neutral" or "contradiction" and nothing else.
Premise and Hypothesis: “‘ {premise} {hypothesis} “‘
Label:
Table 12: The prompts utilized for the e-SNLI dataset.
@@
NORMAL PROMPT UNBOUND You will be provided a biography which is delimited with triple backticks.
Your task is to identify the most important key words from the text that indicate why is the person in the following short bio described as a {label}.
• The key words should be copied verbatim from the text.
• Write the key words as a pipe-separated (|) list of single words.
• Make sure to only output keyword tokens and do not explain the answer.
• Example output: <word1> | <word2> | <word3>
@@
SHORT PROMPT UNBOUND Identify the most important key words from the text that indicate why is the person in the following short bio described as a {label}.
• Write the key words as a pipe-separated (|) list of single words.
Bio: “‘ {bio} “‘
EXTENDED PROMPT UNBOUND Consider the following prompt:
" Determine the occupation of the person in the following short bio.
Give your answer as a single word from (psychologist, surgeon, nurse, dentist, physician) and nothing else.
Bio: “‘ {bio} “‘
Occupation: "
Identify the most important single key words from the whole prompt (Including instruction and bio) that indicate why is the person in the short bio described as a {label}.
Just write the words as a pipe-separated (|) list of words with spaces and do not explain the answer.
NORMAL PROMPT TOP-VAR / TOP-RATIO You will be provided a biography which is delimited with triple backticks.
Your task is to identify the top {k} most important key words from the text that indicate why is the person in the following short bio described as a {label}.
• The key words should be copied verbatim from the text.
• Write the key words as a pipe-separated (|) list of single words.
• Make sure to only output top {k} keyword tokens and do not explain the answer.
• Example output: <word1> | <word2> | <word3>
Bio: “‘ {bio} “‘
SHORT PROMPT TOP-VAR / TOP-RATIO Identify the top {k} most important key words from the text that indicate why is the person in the following short bio described as a {label}.
• Write the key words as a pipe-separated (|) list of single words and do not explain.
Bio: “‘ {bio} “‘
ATTRIBUTION-BASED TOP-VAR / TOP-RATIO
(Selected later from tokens with the highest attribution scores)
Determine the occupation of the person in the following short bio.
Give your answer as a single word from (psychologist, surgeon, nurse, dentist, physician) and nothing else.
Bio: “‘ {bio} “‘
Occupation: {predicted_label}
CLASSIFICATION Determine the occupation of the person in the following short bio.
Give your answer as a single word from (psychologist, surgeon, nurse, dentist, physician) and nothing else.
Bio: “‘ {bio} “‘
Occupation:
Table 13: The prompts utilized for the MedicalBios dataset.

### v2

@@ Appendix A Appendix
Figure A.1: The distribution of predicted labels across epochs of fine-tuning. Pre-trained off-the-shelf models tend to be heavily biased toward a label in poorly performing datasets.
Figure A.2: Token Importance before and after fine-tuning Llama-2-7B based on the InputXGradient explainability method. The predicted word by the model is shown in green and its attributions to previous words are shown in red. The attributions before fine-tuning are more skewed (Fisher-Pearson coefficient of skewness over all the dataset: 3.37±0.31), and become less skewed after fine-tuning (1.42±0.36).
Figure A.3: Token Importance before and after fine-tuning Llama-2-7B based on the Input×Gradient explainability method. The predicted/true label is shown in green, with its attributions to input words in red. The human rationale for the examples are ["dog", "cat], ["twins"], and ["grazes", "touching"]. The fine-tuned model identified these words solely through training on classification data, without any rationale data.
@@
Table A.1: The details of the models we used in this work.
@@
Learning Rate E-SNLI 7e-06
Learning Rate FEVER 7e-06
Learning Rate MedicalBios 3e-06
Num Epochs 10
@@
Table A.2: The hyperparameters used for fine-tuning the models using LoRA.
@@
PROMPT UNBOUND Premise: “‘ {premise} “‘
@@
Identify the most important words from the text that are most relevant to understanding the relationship between the premise and the hypothesis. Write the words as a pipe-separated (|) list of words with spaces. Do not output any other text, symbols, or explanations.
PROMPT TOP-VAR Premise: “‘ {premise} “‘
@@
Identify the top {k} most important words from the text that are most relevant to understanding the relationship between the premise and the hypothesis.
Write the top {k} words as a pipe-separated (|) list of words with spaces. Do not output any other text, symbols, or explanations.
ATTRIBUTION-BASED TOP-VAR
(Selected later from tokens with the highest attribution scores)
@@
Determine the inference relation between two (short, ordered) texts as entailment, if the premise entails the hypothesis, contradiction, if the hypothesis contradicts the premise, or neutral, if neither entailment nor contradiction hold. Respond with exactly one of the following: "entailment", "contradiction", or "neutral." Do not output any other text, symbols, or explanations—just the label.
@@
CLASSIFICATION Premise and Hypothesis: “‘ {premise} {hypothesis} “‘
Determine the inference relation between two (short, ordered) texts as entailment, if the premise entails the hypothesis, contradiction, if the hypothesis contradicts the premise, or neutral, if neither entailment nor contradiction hold. Respond with exactly one of the following: "entailment", "contradiction", or "neutral." Do not output any other text, symbols, or explanations—just the label.
Label:
Table A.3: The prompts utilized for the e-SNLI dataset.
Method Selection Prompt
PROMPT UNBOUND Claim: {claim}
Evidence: “‘ {evidence} “‘
@@
Identify the most important words from the given evidence that are most essential for verifying the factual relationship between the claim and the evidence. Provide these words as a pipe-separated (|) list. Do not output any other text, symbols, or explanations.
PROMPT TOP-VAR Claim: {claim}
Evidence: “‘ {evidence} “‘
Label: {label}
Identify the top {k} most important words from the given evidence that are most essential for verifying the factual relationship between the claim and the evidence. Provide these top {k} words as a pipe-separated (|) list. Do not output any other text, symbols, or explanations.
ATTRIBUTION-BASED TOP-VAR
@@
Claim: {claim}
Evidence: “‘ {evidence} “‘
Determine the factual relationship between the claim and the supporting evidence as "supports" if the evidence supports the claim or "refutes" if the evidence refutes the claim. Respond with exactly one of the following: "supports" or "refutes". Do not output any other text, symbols, or explanations—just the label. Label: {label}
CLASSIFICATION Claim: {claim}
Evidence: “‘ {evidence} “‘
Determine the factual relationship between the claim and the supporting evidence as "supports" if the evidence supports the claim or "refutes" if the evidence refutes the claim. Respond with exactly one of the following: "supports" or "refutes". Do not output any other text, symbols, or explanations—just the label. Label:
Table A.4: The prompts utilized for the FEVER dataset.
@@
PROMPT UNBOUND Bio: “‘ {bio} “‘
Identify the most important words from the text that indicate why is the person in the following short bio described as a {label}. Write the key words as a pipe-separated (|) list of single words. Do not output any other text, symbols, or explanations.
PROMPT TOP-VAR Bio: “‘ {bio} “‘
Identify the top {k} most important words from the text that indicate why is the person in the following short bio described as a {label}. Write the key words as a pipe-separated (|) list of single words. Do not output any other text, symbols, or explanations.
ATTRIBUTION-BASED TOP-VAR
(Selected later from tokens with the highest attribution scores)
@@
Determine the occupation of the person in the short bio. Respond with exactly one of the following: (psychologist, surgeon, nurse, dentist, physician). Do not output any other text, symbols, or explanations—just the label. Occupation: {label}
CLASSIFICATION Bio: “‘ {bio} “‘
Determine the occupation of the person in the short bio. Respond with exactly one of the following: (psychologist, surgeon, nurse, dentist, physician). Do not output any other text, symbols, or explanations—just the label. Occupation:
Table A.5: The prompts utilized for the MedicalBios dataset.